Solve the initial value problem 2x+y-bxy+4y=0 Y()=0 ; yC1=252
Answer #1

de -62 +4 dan 2D - 2D-GP HY :O yet) you: geetest 2(2-2) (2 2a 22y- Gay-uy: 0 ped x = et s) t = logr. x dy = Dy e che zdy,y..) : 212 (2+52)6, +(2-6)(= 252 : solve both equation - (2-12) (cit C2 =0)-) 12-5-26 - 12-82 (2+2) C, +(2-8274 7202) Q+52)6